# xDeepServe: Model-as-a-Service on Huawei CloudMatrix384

### v3

@@ Abstract
Abstract. The rise of scaled-out LLMs and scaled-up SuperPods signals a new era in large-scale AI infrastructure. LLMs continue to scale out via MoE, as seen in recent models like DeepSeek, Kimi, and Qwen. In parallel, AI hardware is scaling up, with Huawei’s CloudMatrix384 SuperPod offering hundreds of GB/s high-speed interconnects. Running large MoE models on SuperPod-scale hardware brings new challenges. It requires new execution models, scalable scheduling, efficient expert load balancing, and elimination of single points of failure.
@@
Figure 1: xDeepServe Architecture over CloudMatrix384 SuperPod.
@@ 1 Introduction
We implement two forms of disaggregation.
@@ 2.2 CloudMatrix384
Figure 2 shows an overview of CloudMatrix384. A CloudMatrix384 SuperPod is a single scale-up domain composed of 48 servers, delivering hundreds of PFLOPs of FP16 compute, several terabytes of on-chip memory, and terabytes per second of memory bandwidth. Each server in the SuperPod is equipped with multiple CPUs, multiple NICs, and 8 Ascend NPU chips. Crucially, the SuperPod integrates three types of network fabrics. The first is a traditional VPC network that connects servers to external systems or cloud services. The second is a scaled-out RoCE network that interconnects all NPU chips and can be extended across multiple SuperPods and previous generation Ascend NPU servers. The third is a scaled-up UB network that makes CloudMatrix384 a SuperPod. It connects all CPUs and NPUs within the SuperPod in an all-to-all fashion, offering roughly several times the bandwidth of RoCE.
@@ 2.3 NPU Execution Mode
  • Graph Mode. Graph mode traces and compiles a static computation graph, which is dispatched as a single unit to the NPU. This removes per-operator launch overhead and maximizes device utilization. Huawei’s TorchAir (built on PTA) supports graph-mode inference on Ascend NPUs. As shown in Figure 3(b), the CPU can dispatch the entire graph to the NPU in one kernel launch. We use graph mode during decode for optimal performance, as input shapes are small and mostly static.
@@ 3 Communication Library over CloudMatrix’s Distributed Shared Memory
A high-performance communication library is essential to fully harness the scaled-up UB fabric in CloudMatrix384. We introduce XCCL, a communication library purpose-built for LLM serving on SuperPod. XCCL proposes distributed memory protocols over CloudMatrix384’s global shared memory to construct efficient networking primitives. Its protocol design resembles classical far-memory systems based on one-siden RDMA verbs farm-nsdi14 ; clover-atc20 . Note that the memory transfer units in NPU’s computing cores, along with the DMA engines, are fundamental to enabling global shared memory in CloudMatrix384. Both can read from or write to the on-chip memory of any other NPU in the SuperPod via the UB fabric. This capability underpins our customized communication library for serving. In general, memory transfer units are used for low-latency communication, while DMA engines are preferred for high-throughput transfers. This is because memory transfer units provide a memory-semantic API and move data between the computing core’s data buffer and on-chip memory, with transfers limited to the buffer size of hundreds of KBs. In contrast, DMA engines use DMA semantics and support bulk data movement up to several GBs.
We describe the point-to-point APIs used in disaggregated prefill and decode (§3.1), the dispatch and combine APIs for expert parallelism (EP) (§3.2), and the A2E/E2A APIs for disaggregated MoE and attention (§3.3).
@@ 3.1 Point-to-Point Primitives
Figure 6: Evaluation of Dispatch/Combine. We vary the batch size per die with a fixed EP128. Dispatch has quantization step.
@@ 3.2 All-to-All Communication for Colocated MoE-Attention
Distributed Memory Protocol of dispatch. The dispatch protocol consists of two main phases: broadcasting the number of tokens each rank should receive, followed by the actual data transfer. Figure 7 illustrates the complete workflow. Step 1: The sender’s serving engine invokes dispatch, passing the source buffer in the app data area, an event ID, and additional parameters. A kernel is launched on the sender NPU, which uses mtu-in to copy data from the app data area to each computing core’s data buffer in parallel. Step 2: If quantization is enabled, the kernel converts the data from FP16/BF16 to INT8 using vector instructions. Step 3: The kernel writes token data into the managed data area, partitioned by destination rank ID. Step 4: The kernel updates each destination rank’s metadata with the number of tokens it will receive. Step 5: Each NPU polls its local metadata area until metadata from all ranks is received. Steps 6–7: Once all metadata is received, the kernel pulls token data from peer NPUs into the data buffer using mtu-in, guided by received offsets and token counts. It then copies the data into the destination buffer in the app data area via mtu-out.
@@ 4.6 Multi-Token Prediction
DeepSeek’s Multi-Token Prediction (MTP) liu2024deepseek-v3 extends conventional next-token prediction by training the model to generate multiple future tokens in sequence. During inference, MTP enables speculative decoding by predicting several tokens in one pass and verifying them afterward, accelerating autoregressive generation.
This section describes how FlowServe integrates MTP into its disaggregated prefill-decode workflow (§5) to improve decode efficiency. Figure 13 illustrates the full execution pipeline.
@@ 9 Contributors
Ao Xiao, Bangzheng He, Baoquan Zhang, Baoxing Huai, Bingji Wang, Bo Wang, Bo Xu, Boyi Hou, Chan Yang, Changhong Liu, Cheng Cui, Chenyu Zhu, Cong Feng, Daohui Wang, Dayun Lin, Duo Zhao, Fengshao Zou, Fu Wang, Gangqiang Zhang, Gengyuan Dan, Guanjie Chen, Guodong Guan, Guodong Yang, Haifeng Li, Haipei Zhu, Hao Feng, Hao Huang, Hao Xu, Hengrui Ma, Hengtao Fan, Hui Liu, Jia Li, Jiang Liu, Jiang Xu, Jie Meng, Jinhan Xin, Junhao Hu, Juwei Chen, Lan Yu, Lanxin Miao, Liang Liu, Linan Jing, Lu Zhou, Meina Han, Mingkun Deng, Mingyu Deng, Naitian Deng, Nizhong Lin, Peihan Zhao, Peng Pan, Pengfei Shen, Ping Li, Qi Zhang, Qin Zhang, Qingrong Xia, Qingyi Zhang, Qunchao Fu, Ren Guo, Ruimin Gao, Shaochun Li, Sheng Long, Shentian Li, Shining Wan, Shuai Shen, Shuangfu Zeng, Shuming Jing, Siqi Yang, Song Zhang, Tao Xu, Tianlin Du, Ting Chen, Wanxu Wu, Wei Jiang, Weinan Tong, Weiwei Chen, Wen Peng, Wenli Zhou, Wenquan Yang, Wenxin Liang, Xiang Liu, Xiaoli Zhou, Xin Jin, Xinyu Duan, Xu Li, Xu Zhang, Xusheng Chen, Yalong Shan, Yang Gan, Yao Lu, Yi Deng, Yi Zheng, Yingfei Zheng, Yiyun Zheng, Yizhou Shan, Yong Gao, Yongqiang Yang, Yuanjin Gong, Yue Yu, Yuetao Chen, Yukun Zhu, Yulong He, Yusu Zhao, Yuyan Wu, Zenan Zhang, Zhaojin Zhuo, Zhaoyang Ji, Zhefeng Wang, Zheng Wang, Zhenhua Yang, Zhenli Sheng, Zhibin Yu, Zhigang Ji, Zhihao Ren, Zhipeng Bian, Zhixia Liu, Zhiyu Dong, Zhonghua Li, Zhou Yu, Zhuoming Shen, Zhuwei Peng, Zi Ye, Zihao Xiang, Zimin Fu, Zixuan Zhang.

### v5

@@ Abstract
The rise of scaled-out LLMs and scaled-up SuperPods signals a new era in large-scale AI infrastructure. LLMs continue to scale out via MoE, as seen in recent models like DeepSeek, Kimi, and Qwen. In parallel, AI hardware is scaling up, with Huawei’s CloudMatrix384 SuperPod offering hundreds of GB/s high-speed interconnects. Running large MoE models on SuperPod-scale hardware brings new challenges. It requires new execution models, scalable scheduling, efficient expert load balancing, and elimination of single points of failure.
@@
Figure 1: xDeepServe over CloudMatrix384 SuperPod. We deploy both disaggregated Prefill–Decode and disaggregated MoE–Attention across heterogeneous clusters. Red lines indicate requests sent to prefill, blue lines represent requests forwarded from prefill to decode.
@@ 1 Introduction
We integrate two forms of disaggregation.
@@ 2.2 CloudMatrix384
Figure 2 shows an overview of CloudMatrix384. A CloudMatrix384 SuperPod is a single scale-up domain composed of 48 servers, delivering hundreds of PFLOPs of FP16 compute, several terabytes of on-chip memory, and terabytes per second of memory bandwidth. Each server in the SuperPod is equipped with multiple CPUs, multiple NICs, and 8 Ascend NPU chips. Crucially, the SuperPod integrates three types of network fabrics. The first is a traditional VPC network that connects servers to external systems or cloud services. The second is a scaled-out RoCE network that interconnects all NPU chips and can be extended across multiple SuperPods and previous generation Ascend NPU servers. The third is a scaled-up UB network that makes CloudMatrix384 a SuperPod. It connects all CPUs and NPUs within the SuperPod in an all-to-all fashion, offering roughly several times the bandwidth of RoCE. See  ascend-npu-hotchips19 ; ascend-c-programming-manual ; cloudmatrix384 for more.
@@ 2.3 NPU Execution Mode
Graph Mode. Graph mode traces and compiles a static computation graph, which is then dispatched to the NPU in a single kernel launch. This eliminates per-operator launch overhead and improves device utilization. Huawei’s TorchAir (based on PTA) enables graph-mode inference on Ascend NPUs. As shown in Figure 3(b), the CPU submits the entire computation graph in one operation, reducing dispatch overhead. We use this mode during decode, where inputs are small and mostly static, to achieve optimal performance.
@@ 3 Communication Library over CloudMatrix’s Distributed Shared Memory
A high-performance communication library is essential to fully harness the scaled-up UB fabric in CloudMatrix384. We introduce XCCL, a communication library purpose-built for LLM serving on SuperPod. XCCL proposes distributed memory protocols over CloudMatrix384’s global shared memory to construct efficient networking primitives. Its protocol design resembles classical far-memory systems based on one-siden RDMA verbs farm-nsdi14 ; clover-atc20 . Note that the memory transfer units (mtu) in NPU’s computing cores, along with the DMA engines, are fundamental to enabling global shared memory in CloudMatrix384. Both can read from or write to the on-chip memory of any other NPU in the SuperPod via the UB fabric. This capability underpins our customized communication library for serving. In general, memory transfer units are used for low-latency communication, while DMA engines are preferred for high-throughput transfers. This is because memory transfer units provide a memory-semantic API and move data between the computing core’s data buffer and on-chip memory, with transfers limited to the buffer size of hundreds of KBs. In contrast, DMA engines use DMA semantics and support bulk data movement up to several GBs.
Below, we describe the point-to-point APIs used in disaggregated prefill and decode (§3.1), the dispatch and combine APIs for expert parallelism (EP) (§3.2), and the A2E/E2A APIs for disaggregated MoE and attention (§3.3).
@@ 3.1 Point-to-Point Primitives
Figure 6: Evaluation of Dispatch/Combine. We vary the batch size per die with a fixed EP128. We use DeepSeek-R1.
@@ 3.2 All-to-All Communication for Colocated MoE-Attention
Distributed Memory Protocol of dispatch. The dispatch protocol consists of two main phases: broadcasting the number of tokens each rank should receive, followed by the actual data transfer. Figure 7 illustrates the complete workflow. Step 1: The sender’s serving engine invokes dispatch, passing the source buffer in the app data area, an event ID, and additional parameters. A kernel is launched on the sender NPU, which uses mtu-in to copy data from the app data area to each computing core’s data buffer in parallel. Step 2: If quantization is enabled, the kernel converts the data from FP16/BF16 to INT8 using certain instructions. Step 3: The kernel writes token data into the managed data area, partitioned by destination rank ID. Step 4: The kernel updates each destination rank’s metadata with the number of tokens it will receive. Step 5: Each NPU polls its local metadata area until metadata from all ranks is received. Steps 6–7: Once all metadata is received, the kernel pulls token data from peer NPUs into the data buffer using mtu-in, guided by received offsets and token counts. It then copies the data into the destination buffer in the app data area via mtu-out.
@@ 4.6 Multi-Token Prediction
This section describes how FlowServe integrates MTP into its disaggregated prefill-decode workflow (§5) to improve decode efficiency. Figure 13 illustrates the full execution pipeline. DeepSeek’s Multi-Token Prediction (MTP) liu2024deepseek-v3 extends conventional next-token prediction by training the model to generate multiple future tokens in sequence. During inference, MTP enables speculative decoding by predicting several tokens in one pass and verifying them afterward, accelerating autoregressive generation.
@@ 9 Contributors
Ao Xiao, Bangzheng He, Baoquan Zhang, Baoxing Huai, Bingji Wang, Bo Wang, Bo Xu, Boyi Hou, Chan Yang, Changhong Liu, Cheng Cui, Chenyu Zhu, Cong Feng, Daohui Wang, Dayun Lin, Duo Zhao, Fengshao Zou, Fu Wang, Gangqiang Zhang, Gengyuan Dan, Guanjie Chen, Guodong Guan, Guodong Yang, Haifeng Li, Haipei Zhu, Haley Li, Hao Feng, Hao Huang, Hao Xu, Hengrui Ma, Hengtao Fan, Hui Liu, Jia Li, Jiang Liu, Jiang Xu, Jie Meng, Jinhan Xin, Junhao Hu, Juwei Chen, Lan Yu, Lanxin Miao, Liang Liu, Linan Jing, Lu Zhou, Meina Han, Mingkun Deng, Mingyu Deng, Naitian Deng, Nizhong Lin, Peihan Zhao, Peng Pan, Pengfei Shen, Ping Li, Qi Zhang, Qian Wang, Qin ZhC Qingrong Xia, Qingyi Zhang, Qunchao Fu, Ren Guo, Ruimin Gao, Shaochun Li, Sheng Long, Shentian Li, Shining Wan, Shuai Shen, Shuangfu Zeng, Shuming Jing, Siqi Yang, Song Zhang, Tao Xu, Tianlin Du, Ting Chen, Wanxu Wu, Wei Jiang, Weinan Tong, Weiwei Chen, Wen Peng, Wenli Zhou, Wenquan Yang, Wenxin Liang, Xiang Liu, Xiaoli Zhou, Xin Jin, Xinyu Duan, Xu Li, Xu Zhang, Xusheng Chen, Yalong Shan, Yang Gan, Yao Lu, Yi Deng, Yi Zheng, Ying Xiong, Yingfei Zheng, Yiyun Zheng, Yizhou Shan, Yong Gao, Yong Zhang, Yongqiang Yang, Yuanjin Gong, Yue Yu, Yuetao Chen, Yukun Zhu, Yulong He, Yusu Zhao, Yuyan Wu, Zenan Zhang, Zhaojin Zhuo, Zhaoyang Ji, Zhefeng Wang, Zheng Wang, Zhenan Fan, Zhenhua Yang, Zhenli Sheng, Zhibin Yu, Zhigang Ji, Zhihao Ren, Zhipeng Bian, Zhixia Liu, Zhiyu Dong, Zhonghua Li, Zhou Yu, Zhuoming Shen, Zhuwei Peng, Zi Ye, Zihao Xiang, Zimin Fu, Zixuan Zhang.